Prove that R^5/R=R⁴ as rings.
1
Expert's answer
2020-03-12T12:00:03-0400

Consider the mapping "f\colon\mathbb R^5\to\mathbb R^4", where "f(a,b,c,d,e)=(a,b,c,d)". It is a ring homomorpism, because "f((a_1,b_1,c_1,d_1,e_1)+(a_1,b_1,c_1,d_1,e_1))="

"=f(a_1+a_2,b_1+b_2,c_1+c_2,d_1+d_2,e_1+e_2)="

"=(a_1+a_2,b_1+b_2,c_1+c_2,d_1+d_2)="

"=(a_1,b_1,c_1,d_1)+(a_2,b_2,c_2,d_2)="

"=f(a_1,b_1,c_1,d_1,e_1)+(a_2,b_2,c_2,d_2,e_2)" and

"f((a_1,b_1,c_1,d_1,e_1)\cdot(a_1,b_1,c_1,d_1,e_1))="

"=f(a_1a_2,b_1b_2,c_1c_2,d_1d_2,e_1e_2)="

"=(a_1a_2,b_1b_2,c_1c_2,d_1d_2)="

"=(a_1,b_1,c_1,d_1)\cdot(a_2,b_2,c_2,d_2)="

"=f(a_1,b_1,c_1,d_1,e_1)\cdot(a_2,b_2,c_2,d_2,e_2)"

So "\mathbb R^5/\ker f\approx Im f"

"Im f=\mathbb R^4" , because for every "(a,b,c,d)\in\mathbb R^4" we have "(a,b,c,d)=f(a,b,c,d,0)" .

Find "\ker f" : if "(a,b,c,d,e)\in\ker f", then "(0,0,0,0)=f(a,b,c,d,e)=(a,b,c,d)" , that is "a=b=c=d=0" , so "\ker f=\{0\}^4\times\mathbb R\approx\mathbb R"

We obtain "\mathbb R^5/\mathbb R\approx\mathbb R^4"